# Imaging the Femtoscopic Freeze-out Source in Heavy Ion Collisions

### v1

@@ Abstract
The impact of the bin-by-bin variation of the SF on the CF has been analyzed. It is observed that variations in the first few bins of the source function (except for the first bin which is below the low limit cut in LL model calculation) significantly impact the CF, whereas the variations at larger $r$ values are less influential. Inversely, the experimental uncertainty of the CFs near the peak ($q\approx 40~{}\rm MeV$) brings considerable variation to the SF. In order to achieve more accurate imaging, it is demanded to further reduce the uncertainty of the CF, particularly at small $q$. If given high-accuracy experimental CFs, the direct imaging technique developed here offers a novel opportunity to probe the proton density profile of the exotic bubble nucleus, which has been recently proposed [41] but yet in debate [42], in the relativistic collisions induced by those nuclei on future radioactive ion beam facilities. Nevertheless, due to the non-negativity constraint required by the Bayesian-based deblurring algorithm, the method is limited by the choice of basis functions for representing the source. Future investigations are required to explore new inversion techniques such as neural networks, and to adopt mathematically more suitable source bases, as well as to optimize the selection of the convergence criteria, which will all advance the development and application of imaging techniques based on correlation function method.
@@
To summarize, a direct imaging method on femtometer scale has been realized by combining the Richardson-Lucy deblurring algorithm and the intensity interferometry in ultra relativistic heavy ion collisions. This method allows for a more accurate extraction of source distributions and final state interaction between particle pair based on the Lednický-Lyuboshitz model analysis. Applying this imaging method to $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au at center-of-mass energy 200 GeV measured by the STAR collaboration, we observe the same freeze-out source functions for $p$ and $\bar{p}$ emissions with distinct non-Gaussian characteristics. These findings confirm the matter-antimatter symmetry in coordinate space once produced and suggest the ultrafast feature of the evolution of the collision, which is hardly accessible through other experimental means alone. The new method can be extended to three dimensional imaging. It represents a significant advancement of our ability to image the femtoscopic heavy ion collisions and to explore matter-antimatter physics in terrestrial laboratories.

### v2

@@ IV Application to data
The impact of the bin-by-bin variation of the SF on the CF has been analyzed. It is observed that variations in the first few bins of the source function (except for the first bin which is below the low limit cut in LL model calculation) significantly impact the CF, whereas the variations at larger $r$ values are less influential. Inversely, the experimental uncertainty of the CFs near the peak ($q\approx 40~{}\rm MeV$) brings considerable variation to the SF. In order to achieve more accurate imaging, it is demanded to further reduce the uncertainty of the CF, particularly at small $q$. If given high-accuracy experimental CFs, the technique developed here offers a novel opportunity to identify exotic nuclear structure on future radioactive ion beam facilities. For instance, one can probe the proton density profile of bubble nucleus which has been recently proposed [41] but yet in debate [42]. Moreover, the imaging method developed in this work is compatible with machine learning techniques, offering a promising avenue for further enhancement and application in nuclear physics research.
V Summary
@@
A direct imaging method on femtometer scale has been realized by combining the Richardson-Lucy deblurring algorithm and the intensity interferometry in ultra relativistic heavy ion collisions. This method allows for a more accurate extraction of source distributions and final state interaction between particle pair based on the Lednický-Lyuboshitz model analysis. Applying this imaging method to $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au at center-of-mass energy 200 GeV measured by the STAR collaboration, we observe the same freeze-out source functions for $p$ and $\bar{p}$ emissions with distinct non-Gaussian characteristics.These findings confirm the matter-antimatter symmetry in coordinate space once produced and suggest the ultrafast feature of the evolution of the collision, which is hardly accessible through other experimental means alone. The new method can be extended to three dimensional imaging and is inherently compatible with machine learning approaches. It represents a significant advancement of our ability to image the femtoscopic heavy ion collisions and to explore matter-antimatter physics in terrestrial laboratories.